# Dyadic Interaction Modeling for Social Behavior Generation

### v2

@@ Abstract
Human-human communication is like a delicate dance where listeners and speakers concurrently interact to maintain conversational dynamics. Hence, an effective model for generating listener nonverbal behaviors requires understanding the dyadic context and interaction. In this paper, we present an effective framework for creating 3D facial motions in dyadic interactions. Existing work consider a listener as a reactive agent with reflexive behaviors to the speaker’s voice and facial motions. The heart of our framework is Dyadic Interaction Modeling (DIM), a pre-training approach that jointly models speakers’ and listeners’ motions through masking and contrastive learning to learn representations that capture the dyadic context. To enable the generation of non-deterministic behaviors, we encode both listener and speaker motions into discrete latent representations, through VQ-VAE. The pre-trained model is further fine-tuned for motion generation. Extensive experiments demonstrate the superiority of our framework in generating listener motions, establishing a new state-of-the-art according to the quantitative measures capturing the diversity and realism of generated motions. Qualitative results demonstrate the superior capabilities of the proposed approach in generating diverse and realistic expressions, eye blinks and head gestures.
@@ 1 Introduction
Despite considerable progress, existing methods are limited, particularly in capturing real-life conversation dynamics and failing to generalize across listeners. Relevant work focused on either generating speakers’ or listeners’ movements. Speaker motion generation [10, 60, 33, 19, 55, 64, 8, 68, 61] primarily focuses on producing videos where lip movements are synchronized with speech, a.k.a., lip syncing. Listener motion generation, on the other hand, automatically interprets the speaker’s voice and nonverbal behaviors to generate the listener’s nonverbal responses. Traditionally, listener motion generation have not fully addressed the intricacies of dyadic interactions, which involve more than simple reactions to the speaker’s utterances including a multifaceted, concurrent exchange of signals and emotions  [78, 47, 62]. They were also sometimes tuned to specific listeners and unable to generalize to others [47]. Earlier work used heuristics and simple machine learning models to generate a limited set of reactions, e.g., head nods [27]. The Responsive Listening Head Generation (RLHG)[78] approach mimics the regression techniques found in Speaker Generation[10], leading to a reduction in the unpredictability of responses and a smoothing over of listener motions. Ng et al.  [47] proposed Learning2Listen (L2L) to mitigate this issue through a motion categorization strategy. In L2L, a codebook is learned through a VQ-VAE [45] that can be used to generate plausible and smooth behaviors. Emotional Listener Portrait (ELP) [62] utilizes an Adaptive Space Encoder to map the discretized features combined with emotion to the listener’s motion parameters. Similarly, Speaker motion generation techniques [55, 21, 71] have struggled to incorporate the feedback loop present in actual dyadic interactions, where a speaker might adjust their speech and expressions in response to the listener’s behavior. These shortcomings underscore a critical gap in the simulation of naturalistic human interactions, highlighting the need for a more holistic and dyadic modeling approach.
@@ 2.1 Speech-driven Speaker Generation
There is a wide body of work on automatic facial motion generation [70, 39, 4, 81, 37, 24, 67, 38]. A large number of studies focus on 2D facial animation [20, 13, 9, 52, 68, 16, 73, 8, 33, 76]. Traditional procedural techniques [43, 66, 72, 18] rely on a framework of explicit rules for simulating the movements of a talking mouth, such as the use of dominance functions [43] for delineating speech control parameters.
These procedural strategies offer precise control over mouth movement accuracy but are labor-intensive due to the extensive manual tuning required. In contrast, data-driven methods [5, 42, 36, 65, 51, 14, 32, 55] present an alternative, proposing various techniques to generate 3D facial animations with less manual intervention. Among these, Karras et al. [36] propose an end-to-end network with linear predictive coding to generate audio and novel latent codes with facial expression variations. In Zhou et al. [80], the proposed network predicts viseme curves by integrating phoneme groups, landmarks, and audio signals. VOCA [14] offers speaker-independent animation capturing diverse speaking styles, though it primarily affects the lower face. More recently, MeshTalk [55] has made strides in disentangling audio-correlated and uncorrelated facial motions through a categorical latent space. Innovations continue with FaceFormer [21] and CodeTalker [71], which leverage long-term audio context via transformer-based models and augment them with a VQ-VAE motion prior for autoregressive motion synthesis, respectively.
@@ 3.2 Learning discrete Speaker-Listener motions
Formally, we train two separate VQ-VAE models, denoted as $VQ^{(s)}$ and $VQ^{(l)}$, to handle speaker and listener motions, respectively. For clarity, we present only the broad overview of the VQ-VAE workflow for encoding and reconstructing listener motions. In particular, the encoder $Enc^{(l)}_{VQ}$ processes the listener motions $L$ to produce the continuous latent representation $z^{cont}_{L}=Enc^{(l)}_{VQ}(L)$. The latent representation is then quantized to $z^{(l)}$ using the nearest vector from the shared codebook $C$. The reconstruction of the listener motion is then given by: $\hat{L}=Enc^{(l)}_{VQ}(z^{(l)}).$
@@
$\mathcal{L}_{codebook}^{(l)}=\|sg[E(x)]-q\|_{2}^{2}+\beta\|sg[q]-E(x)\|_{2}^{2},$ (2)
@@ 3.3 Dyadic Interaction Modeling
The overall pipeline is illustrated in Figure 2. Given $s_{1:T}$ and $l_{1:T}$, DIM-Listener first generates $z^{(s)}$ and $z^{(l)}$ as the discrete motion sequences produced by the learned motion codebooks: $z^{(s)}=Enc^{(s)}_{VQ}(s_{1:T})$ and $z^{(l)}=Enc^{(l)}_{VQ}(l_{1:T})$, and uses these sequences as the discrete latent codes to later train the network. It is important to note that both $Enc^{(s)}_{VQ}$ and $Enc^{(l)}_{VQ}$ are frozen to produce consistent discrete latent codes. Then, we uniformly mask $p\%$ of frames for both speakers and listeners, i.e.,
$s_{masked}=Mask_{p}(s+E_{s}+E^{P}_{s})\;;\;z^{(s)}_{masked}=Mask_{p}(z^{(s)})$ (4)
$l_{masked}=Mask_{p}(l+E_{l}+E^{P}_{l})\;;\;z^{(l)}_{masked}=Mask_{p}(z^{(l)})$
where $E_{s}$ and $E_{l}$ denote the learnable speaker and listener embeddings and $E^{P}_{s}$ and $E^{P}_{l}$ denote the sinusoidal positional embeddings.
@@
$\mathcal{L}_{c}=-\frac{1}{N}\Sigma_{i=1}^{N}log[\frac{exp(x^{(s)}_{i}\cdot x^{(l)}_{i}/\tau)}{\Sigma{i\neq k}exp(x^{(s)}_{i}\cdot x^{(l)}_{k}/\tau)+exp(x^{(s)}_{i}\cdot x^{(l)}_{i}/\tau)}]$ (5)
@@
We further concatenated $x^{(s)}_{masked}$ and $x^{(l)}_{masked}$ along the feature dimension to form $x^{(s-l)}_{masked}$ and forward the fused feature into a joint Transformer encoder to produce $j^{(s-l)}_{masked}$ and later split $j^{(s-l)}_{masked}$ back into separate speaker and listener features, i.e.,
@@
The predicted discrete motions $\tilde{x^{(s)}}$ and $\tilde{x^{(l)}}$ are then forwarded into $Dec^{(s)}_{VQ}$ and $Dec^{(l)}_{VQ}$ for final continuous motion predictions.
$\tilde{s}=Dec^{(s)}_{VQ}z^{\prime(s)})\ \;;\;\tilde{l}=Dec^{(l)}_{VQ}(z^{\prime(l)})$ (8)
@@
$\mathcal{L}^{(s)}_{rec}=\Sigma_{t\in M}-logp(z^{(s)}_{t}|z^{\prime(s)}_{t})+||s_{t}-\tilde{s}_{t}||_{2}$ (9)
$\mathcal{L}^{(l)}_{rec}=\Sigma_{t\in M}-logp(z^{(l)}_{t}|z^{\prime(l)}_{t})+||l_{t}-\tilde{l}_{t}||_{2}$
@@ 4.1 Experimental Settings
We perform the Dyadic Interaction Modeling on CANDOR [53] dataset. It consists of a collection of 1,656 conversations in English. With a total duration of over 850 hours, this large corpus includes more than 7 million words, encompassing rich audio, video, and speech transcripts. It enables detailed moment-to-moment analysis of vocal and facial expressions as well as semantic and conversational contexts.
Implementation Details For the pre-training stage, we use an Adam optimizer with a learning rate of $1e^{-5}$ and train the model on CANDOR [53] for $100$ epochs. We use the same architectures for the Transformer encoders and decoders of DIM-Listener and VQ-VAE, which consists of a transformer with $8$ hidden layers, $8$ attention heads and an intermediate size of $768$. For pre-processing of CANDOR, we use the provided transcript with the dataset to extract utterance-level samples of speaker-listener behaviors. For pre-processing of ViCo [2] and LM_Listener [48], we follow the pre-processing guidelines and train-test splits from prior works [47, 2, 62] for fair comparisons. For all experiments, we extract EMOCA [15], an expressive 3DMMs parameter set, from visual data and adopt it as facial representation for all methods and re-train them for the sake of fair comparison. We follow previous work [77, 47, 62] and adopt expression code (Exp) and pose code (Pose) of EMOCA [15] parameters as the description of the speaker and listener’s motions. We use HuBERT [31] to extract audio features. Since HuBERT is already a robust pre-trained speech encoder, we do not use the audio features in our encoders and directly feed them into our model’s decoder. For the implementation source code of our method, please see supplementary materials.
@@
  • Random: We randomly select the facial and head motion parameters in the training data, and inject random small perturbations into the normal distribution.
  • Mirror: We use the smoothed speaker’s motion as the listener’s motion.
  • NN-motion/NN-audio: For arbitrary input speaker motion or audio input, we find its nearest neighbor from the training set and use its corresponding listener motion as output.
Metrics We re-implement all baselines and train them with EMOCA 3DMM as output motion representation. We then evaluate them using the following metrics reported in LM_Listener [48]:
More details about evaluation metrics are available in the supplementary materials.
@@ 4.2 Benchmark Performances
Qualitative Results. We first compare with the photorealistic results of RLHG [78] and L2L [47] in Figure 4. We feed the EMOCA [15] 3DMM predictions and the listener appearance into PIRenderer [54] and render the corresponding images frame by frame. Since there is no guideline or source code on the renderer’s settings in L2L [47], we use PIRenderer [69] official code for training and generation. We fine-tune a unique renderer for each speaker-listener pair and render the listener image sequence using the same fine-tuned weight of each pair for all approaches. The listener head synthesized by RLHG [78]. L2L [47], and ELP [62] exhibit several shortcomings compared to ours: 1) our method demonstrates more expressive behaviors that synchronize with speakers’ and target listeners’ sentiments (first example), 2) our method can generate diverse head movements compared to other methods (second example). We also demonstrate our method’s ability to model emotions and facial behaviors of our model in Figure 5. In particular, we show that DIM-Listener can generate multiple types of social signals such as smiles, head shaking and blinks. These findings suggest the benefits of Dyadic Interaction Modeling of DIM-Listener on the large-scale CANDOR [53] dataset with more diverse dyadic behaviors. We’ll discuss the limitations of our framework in the supplementary materials.
@@ 4.3 Ablation Analysis
that fixes the VQ-Decoder to decode continuous signals from discrete motion predictions, integrating the VQ-Decoder into the listener motion prediction model notably enhances performance, ranking as the second most important component in DIM-Listener. While $\mathcal{L}_{c}$ ranks as the least important component, it still contributes significantly to performance by encouraging role-specific encoders to extract meaningful information from input signals relevant not only to the current person but also the conversational partner. Finally, we demonstrate the benefits of jointly modeling both speaker and listener behaviors with $S-L$, which aligns with existing literature indicating that speaker and listener behaviors correlate and mutually influence each other.
@@ 5 Conclusion
In this work, we propose Dyadic Interaction Modeling, a self-supervised pre-training strategy that enhances the model’s ability to encode a unified representation from both speaker and listener motions through the utilization of self-supervised contrastive learning, focusing on the reconstruction of masked-hidden units. Leveraging DIM, we develop DIM-Listener for listener motion generation in dyadic interactions with detailed and realistic facial expressions and head motions. Thanks to DIM, our method is not only able to generate listener behaviors from speaker audio-visual input but also able to generate speaker facial behaviors from speaker speech (DIM-Speaker). Extensive experiments and visualizations on dyadic motion generation demonstrate the effectiveness of our approach.
6 Acknowledgement

### v3

# %TODO␣FINAL:␣Replace␣with␣your␣institution␣list.https://boese0601.github.io/dim/

@@ Abstract
Human-human communication is like a delicate dance where listeners and speakers concurrently interact to maintain conversational dynamics. Hence, an effective model for generating listener nonverbal behaviors requires understanding the dyadic context and interaction. In this paper, we present an effective framework for creating 3D facial motions in dyadic interactions. Existing work consider a listener as a reactive agent with reflexive behaviors to the speaker’s voice and facial motions. The heart of our framework is Dyadic Interaction Modeling (DIM), a pre-training approach that jointly models speakers’ and listeners’ motions through masking and contrastive learning to learn representations that capture the dyadic context. To enable the generation of non-deterministic behaviors, we encode both listener and speaker motions into discrete latent representations, through VQ-VAE. The pre-trained model is further fine-tuned for motion generation. Extensive experiments demonstrate the superiority of our framework in generating listener motions, establishing a new state-of-the-art according to the quantitative measures capturing the diversity and realism of generated motions. Qualitative results demonstrate the superior capabilities of the proposed approach in generating diverse and realistic expressions, eye blinks and head gestures. The code is available at https://github.com/Boese0601/Dyadic-Interaction-Modeling.
@@ 1 Introduction
Despite considerable progress, existing methods are limited, particularly in capturing real-life conversation dynamics, and difficult to generalize across listeners. Relevant work focused on either generating speakers’ or listeners’ movements. Speaker motion generation [11, 64, 34, 20, 59, 70, 9, 74, 65] primarily focuses on producing videos where lip movements are synchronized with speech, a.k.a., lip syncing. Listener motion generation, on the other hand, automatically interprets the speaker’s voice and nonverbal behaviors to generate the listener’s nonverbal responses. Traditionally, listener motion generation have not fully addressed the intricacies of dyadic interactions, which involve more than simple reactions to the speaker’s utterances including a multifaceted, concurrent exchange of signals and emotions  [84, 50, 66]. They were also sometimes tuned to specific listeners and unable to generalize to others [50]. Earlier work used heuristics and simple machine learning models to generate a limited set of reactions, e.g., head nods [28]. The Responsive Listening Head Generation (RLHG)[84] approach mimics the regression techniques found in Speaker Generation[11], leading to a reduction in the unpredictability of responses and a smoothing over of listener motions. Ng et al.  [50] proposed Learning2Listen (L2L) to mitigate this issue through a motion categorization strategy. In L2L, a codebook is learned through a VQ-VAE [48] that can be used to generate plausible and smooth behaviors. Emotional Listener Portrait (ELP) [66] utilizes an Adaptive Space Encoder to map the discretized features combined with emotion to the listener’s motion parameters. Similarly, Speaker motion generation techniques [59, 22, 78] have struggled to incorporate the feedback present in actual dyadic interactions. These shortcomings underscore a critical gap in the simulation of naturalistic human interactions, highlighting the need for a more holistic and dyadic modeling approach.
@@ 2.1 Speech-driven Speaker Generation
There is a wide body of work on automatic facial motion generation [76, 41, 4, 87, 38, 25, 73, 40, 77, 39, 68, 54, 7]. A large number of studies focus on 2D facial animation [21, 14, 10, 56, 74, 17, 80, 9, 34, 83]. Traditional procedural techniques [46, 72, 79, 19] rely on a framework of explicit rules for simulating the movements of a talking mouth, such as the use of dominance functions [46] for delineating speech control parameters.
These procedural strategies offer precise control over mouth movement accuracy but are labor-intensive due to the extensive manual tuning required. In contrast, data-driven methods [5, 45, 37, 71, 55, 15, 33, 59, 69] present an alternative, proposing various techniques to generate 3D facial animations with less manual intervention. Among these, Karras et al. [37] propose an end-to-end network with linear predictive coding to generate audio and novel latent codes with facial expression variations. In Zhou et al. [86], the proposed network predicts viseme curves by integrating phoneme groups, landmarks, and audio signals. VOCA [15] offers speaker-independent animation capturing diverse speaking styles, though it primarily affects the lower face. More recently, MeshTalk [59] has made strides in disentangling audio-correlated and uncorrelated facial motions through a categorical latent space. Innovations continue with FaceFormer [22] and CodeTalker [78], which leverage long-term audio context via transformer-based models and augment them with a VQ-VAE motion prior for autoregressive motion synthesis, respectively.
@@ 3.2 Learning discrete Speaker-Listener motions
Formally, we train two separate VQ-VAE models, denoted as $VQ^{(s)}$ and $VQ^{(l)}$, to handle speaker and listener motions, respectively. For clarity, we present only the broad overview of the VQ-VAE workflow for encoding and reconstructing listener motions. In particular, the encoder $Enc^{(l)}_{VQ}$ processes the listener motions $l_{1:T}$ to produce the continuous latent representation $z^{(l)}_{cont}=Enc^{(l)}_{VQ}(l_{1:T})$. The latent representation is then quantized to $z^{(l)}$ using the nearest vector from the shared codebook $C$. The reconstruction of the listener motion is then given by: $\hat{l}_{1:T}=Dec^{(l)}_{VQ}(z^{(l)}).$
@@
$\mathcal{L}_{codebook}^{(l)}=\|sg[E(l_{1:T})]-z^{(l)}\|_{2}^{2}+\beta\|sg[z^{(l)}]-E(l_{1:T})\|_{2}^{2},$ (2)
@@ 3.3 Dyadic Interaction Modeling
The overall pipeline is illustrated in Figure 2. Given $s_{1:T}$ and $l_{1:T}$, DIM-Listener first generates $z^{(s)}$ and $z^{(l)}$ as the discrete motion sequences produced by the learned motion codebooks: $z^{(s)}=Enc^{(s)}_{VQ}(s_{1:T})$ and $z^{(l)}=Enc^{(l)}_{VQ}(l_{1:T})$, and uses these sequences as the discrete latent codes to later train the network. It is important to note that both $Enc^{(s)}_{VQ}$ and $Enc^{(l)}_{VQ}$ are frozen to produce consistent discrete latent codes. Then, we independently mask $p\%$ of random frames for both speakers and listeners, i.e.,
$s_{masked}=Mask_{p}(s+E_{s}+E^{P}_{s})\;;\;l_{masked}=Mask_{p}(l+E_{l}+E^{P}_{l})$ (4)
where $E_{s}$ and $E_{l}$ denote the learnable speaker and listener embeddings and $E^{P}_{s}$ and $E^{P}_{l}$ denote the sinusoidal positional embeddings. $s$ and $l$ are speaker motions and the listener motions from ground truth.
@@
$\mathcal{L}_{c}=-\frac{1}{N}\Sigma_{i=1}^{N}log[\frac{exp(x^{(s)}_{i}\cdot x^{(l)}_{i}/\tau)}{\Sigma_{i\neq k\;\;}exp(x^{(s)}_{i}\cdot x^{(l)}_{k}/\tau)+exp(x^{(s)}_{i}\cdot x^{(l)}_{i}/\tau)}]$ (5)
@@
We further concatenated $x^{(s)}_{masked}$ and $x^{(l)}_{masked}$ along the feature dimension and forward the fused feature into a joint Transformer encoder to produce $j^{(s-l)}_{masked}$ and later split $j^{(s-l)}_{masked}$ back into separate speaker and listener features, i.e.,
@@
The predicted discrete motions $z^{\prime(s)}$ and $z^{\prime(l)}$ are then forwarded into $Dec^{(s)}_{VQ}$ and $Dec^{(l)}_{VQ}$ for final continuous motion predictions.
$\tilde{s}=Dec^{(s)}_{VQ}(z^{\prime(s)})\ \;;\;\tilde{l}=Dec^{(l)}_{VQ}(z^{\prime(l)})$ (8)
@@
$\mathcal{L}^{(s)}_{rec}=\Sigma_{t\in M}-logp(z^{(s)}_{t}|z^{\prime(s)}_{t})+||s_{t}-\tilde{s}_{t}||_{2}^{2}$ (9)
$\mathcal{L}^{(l)}_{rec}=\Sigma_{t\in M}-logp(z^{(l)}_{t}|z^{\prime(l)}_{t})+||l_{t}-\tilde{l}_{t}||_{2}^{2}$
@@ 4.1 Experimental Settings
Dyadic Interaction Modeling is performed on CANDOR [57] dataset. It consists of a collection of 1,656 conversations in English. With a total duration of over 850 hours, this large corpus includes more than 7 million words, encompassing rich audio, video, and speech transcripts. It enables detailed moment-to-moment analysis of vocal and facial expressions as well as semantic and conversational contexts. CANDOR provides speech diarization and alignment, which was used for speaker/listener designation at the utterance level. The dataset also provides separate speech tracks, so there is no overlapping speech.
Implementation Details For the pre-training stage, we use an Adam optimizer with a learning rate of $1e^{-5}$ and train the model on CANDOR [57] for $100$ epochs. We use a masking ratio $p=75\%$, which is determined with hyper-parameter tuning with respect to the pre-training validation reconstruction loss. We use HuBERT [32] as our audio feature extractor. We use the same architectures for the Transformer encoders and decoders of DIM-Listener and VQ-VAE, which consists of a transformer with $8$ hidden layers, $8$ attention heads and an intermediate size of $768$. For pre-processing of CANDOR, we use the provided transcript with the dataset to extract utterance-level samples of speaker-listener behaviors. For pre-processing of ViCo [2] and LM_Listener [51], we follow the pre-processing guidelines and train-test splits from prior works [50, 2, 66] for fair comparisons. For all experiments, we extract EMOCA [16], an expressive 3DMMs parameter set, from visual data and adopt it as facial representation for all methods and re-train them for the sake of fair comparison. We follow previous work [84, 50, 66] and adopt expression code (Exp) and pose code (Pose) of EMOCA [16] parameters as the description of the speaker and listener’s motions. We use HuBERT [32] to extract audio features. Since HuBERT is already a robust pre-trained speech encoder, we do not use the audio features in our encoders and directly feed them into our model’s decoder. For the implementation source code of our method, please see supplementary materials.
@@
  • We randomly select the facial and head motion parameters in the training data, and inject random small perturbations into the normal distribution.
  • We use the smoothed speaker’s motion as the listener’s motion.
  • For arbitrary input speaker motion or audio input, we find its nearest neighbor from the training set and use its corresponding listener motion as output.
Metrics We re-implement all baselines and train them with EMOCA 3DMM as output motion representation. We then evaluate them using the metrics reported in LM_Listener [51] and ELP [66], namely the Frechet Distance (FD), Paired FD (P-FD) for synchrony , Mean Squared Error (MSE), SI for Diversity (SID), Variance (Var), and Residual Pearson Correlation Coefficient (rPCC). We provide more details about the used metrics in the supplementary materials.
@@ 4.2 Benchmark Performances
Qualitative Results. We first compare with the photorealistic results of RLHG [84] and L2L [50] in Figure 4. We feed the EMOCA [16] 3DMM predictions and the listener appearance into PIRenderer [58] and render the corresponding images frame by frame. Since there is no guideline or source code on the renderer’s settings in L2L [50], we use PIRenderer [75] official code for training and generation. We fine-tune a unique renderer for each speaker-listener pair and render the listener image sequence using the same fine-tuned weight of each pair for all approaches. Compared to RLHG [84], L2L [50], and ELP [66]: 1) our method demonstrates more expressive behaviors that synchronize with speakers’ and target listeners’ sentiments (first example), 2) our method can generate diverse head movements compared to other methods (second example). We also demonstrate our method’s ability to model emotions and facial behaviors of our model in Figure 5. In particular, we show that DIM-Listener can generate multiple types of social signals such as smiles, head shaking and blinks. These suggest the benefits of Dyadic Interaction Modeling of DIM-Listener on the large-scale CANDOR [57] dataset with more diverse dyadic behaviors.
4.3 Ablation Analysis
that fixes the VQ-Decoder to decode continuous signals from discrete motion predictions, integrating the VQ-Decoder into the listener motion prediction model notably enhances performance, ranking as the second most important component in DIM-Listener. While $\mathcal{L}_{c}$ ranks as the least important component, it still contributes significantly to performance by encouraging role-specific encoders to extract meaningful information from input signals relevant not only to the current person but also the conversational partner. Finally, we demonstrate the benefits of jointly modeling both speaker and listener behaviors with $S-L$, which aligns with existing literature indicating that speaker and listener behaviors correlate and mutually influence each other. we also provide modality ablation as well as online listener behavior generation results in the supplementary material.
@@ 5 Conclusion
In this study, we introduced Dyadic Interaction Modeling, a novel self-supervised pre-training strategy designed to improve the model’s capability to effectly encode representations from both speaker and listener motions via large-scale dyadic data. Our comprehensive experiments and detailed visualizations demonstrate the capability of DIM-Listener and DIM-Speaker in generating accurate and realistic motions for both listeners and speakers.
Acknowledgement